OpenStudy (anonymous):

What is the recipricol of -7?

OpenStudy (tkhunny):

Do you have a definition of "reciprocal"? \(\dfrac{1}{-7} = -\dfrac{1}{7}\)

OpenStudy (anonymous):

yes thankyou. what would -1/7 + -2 be?

OpenStudy (tkhunny):

I don't understand that notation. Is it \(-\dfrac{1}{7} + (-2)\)?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

First, resolve the silly notation that requires parentheses: \(-\dfrac{1}{7} - 2\) Second, find a common denominator and add. You do that part.